# The Harmonic Exponential Filter for Nonparametric Estimation on Motion Groups

### v1

@@ Abstract
Bayesian estimation is a vital tool in robotics as it allows systems to update the belief of the robot state using incomplete information from noisy sensors. To render the state estimation problem tractable, many systems assume that the motion and measurement noise, as well as the state distribution, are all unimodal and Gaussian. However, there are numerous scenarios and systems that do not comply with these assumptions. Existing non-parametric filters that are used to model multimodal distributions have drawbacks that limit their ability to represent a diverse set of distributions. In this paper, we introduce a novel approach to nonparametric Bayesian filtering to cope with multimodal distributions using harmonic exponential distributions. This approach leverages two key insights of harmonic exponential distributions: a) the product of two distributions can be expressed as the element-wise addition of their log-likelihood Fourier coefficients, and b) the convolution of two distributions can be efficiently computed as the tensor product of their Fourier coefficients. These observations enable the development of an efficient and exact solution to the Bayes filter up to the band limit of a Fourier transform. We demonstrate our filter’s superior performance compared with established nonparametric filtering methods across a range of simulated and real-world localization tasks.  [^1] Project page https://montrealrobotics.ca/hef/
@@ I Introduction
Recursive Bayesian filtering is a powerful approach for state estimation in robotics. It allows the fusion of several different sensor streams while producing belief estimations over the state that can be used in downstream tasks such as planning, control, and mapping. To render the state estimation problem computationally tractable, certain assumptions are frequently invoked, including the presumption that the state distribution, as well as the motion and measurement noise, are unimodal and Gaussian distributed. However, there are many scenarios where a unimodal Gaussian belief fails to place a high probability (density) over the true state. Estimates involving uncertain data associations [1], propagating errors through rigid-body motions [2], and measurements that suffer from multipath errors [3] and symmetries [4] all result in distributions that are distorted or multimodal.
@@
Beyond Gaussian methods, nonparametric filtering offers an alternative to filters reliant on unimodal distributions or Gaussian assumptions. Examples include particle filters [7, 8, 9] and histogram filters [10, 11, 12]. However, these approaches are not free of limitations. Particle filters are susceptible to particle deprivation [1, Chapter 4], where the correct state is no longer represented within the set of particles. Similarly, histogram filters require resource-intensive computations, including convolutions during the motion propagation phase, and are only accurate up to the quantization error when computing the probability bins.
In this paper, we propose a novel nonparametric and multimodal filtering approach that leverages harmonic exponential distributions [13]. We name this method the harmonic exponential filter (HEF) and it relies on three key benefits of harmonic exponential distributions. First, the prediction step takes advantage of the Fourier analysis convolution theorem, yielding an efficient tensor product instead of a convolution during motion propagation. Second, the update step of the Bayesian filter becomes an element-wise addition. Finally, the Fourier coefficient representation of the harmonic exponential distribution represents more naturally smoothly-varying density functions when compared to histogram binning. These steps are illustrated in Fig 1.
@@
  • An exact approach to computing the posterior belief of the Bayes filter on a compact Lie group, for band-limited prior, motion, and measurement likelihood.
  • A novel nonparametric recursive filtering approach for non-commutative motion groups that can capture multimodal distributions.
  • A filtering strategy whose formulation respects the structure of motion groups and their symmetries.
@@ II Related Work
Various approaches have extended Bayesian filtering beyond the unimodal Gaussian assumptions of the Kalman filter (KF). For instance, the KF has been adapted to operate with a mixture of parameterized distributions instead of a unimodal Gaussian [14]. Gilitschenski et al. [15] adapted a mixture model to use direction statistics on the manifold of rotation groups. However, a key challenge with such approaches is the growth of the number of mixture components with each iteration. Common approaches to overcome this challenge are based on distance metrics to merge mixture components [16] or resampling strategies [17].
The KF has also been extended to exploit symmetries of the system by characterizing the state of the robot using Lie groups [18]. If the state error is modeled in the Lie algebra of motion groups with a Gaussian belief and then mapped back to the group with the exponential map, a non-Gaussian posterior is obtained on the manifold [5, 2]. However, to propagate the error state over time and attain a tractable solution, these methods rely on second-order approximations.
@@
Another family of nonparametric and multimodal filters are particle filters which represent the posterior belief using a discrete set of particles [7]. These methods can model general distributions, but are known to suffer from particle deprivation, particularly in high-dimensional state spaces [1, Chapter 4]. Research efforts have been devoted to addressing this problem by proposing better sampling schemes [8, 9], and by changing how particles evolve in the prediction step to better match the posterior distribution[20]. In contrast, our method does not use particles to model the posterior belief and does not suffer from particle deprivation.
Kernel density estimates are a prevalent method to represent nonparametric probability density functions. They can be adapted to handle belief propagation through probabilistic graphical models [21], and the product of distributions within a filtering framework [22].
Pfaff et al. [23] presented an approach that is conceptually similar to ours that they called Fourier filtering. They showed their work on the unit circle and the unit sphere [24, 25]. Although their approach resembles ours in that the probability density functions are represented by Fourier coefficients, our approach extends this concept in several meaningful ways. First, since we model density functions in the log-likelihood space, the product of distributions becomes an element-wise addition instead of the more intricate tensor product needed for Fourier filtering. Second, by representing distributions with the Fourier coefficients derived from the log-likelihood function, the corresponding probability density function is guaranteed to be non-negative. Lastly, our approach can represent densities on locally compact subsets of motion groups $SE(N)$, particularly, any matrix Lie group and not just on a circle or a sphere.
@@ III Harmonic Exponential Filtering
Fourier analysis provides a way to parameterize well-behaved functions in $\mathbb{R}^{N}$ (i.e. one that is square integrable $\mathcal{L}^{2}$ on $\mathcal{V}\subseteq\mathbb{R}^{N}$ and has a finite number of discontinuities in any finite interval in $\mathbb{R}^{N}$) using an infinite sum of trigonometric basis functions. This approach has proven to be useful for various engineering applications for several reasons. Moreover, the convolution theorem shows that computing integrals in spatial domains can be replaced by an efficient tensor product in the frequency domain. A generalized form of Fourier analysis can be used on locally compact subsets of motion groups. The topic is presented in depth by Chirikjian and Kyatkin in [26, Chapters 8,10]. This section presents an overview of our approach. First by providing a brief introduction to Fourier analysis on motion groups (Sec. III-A), then showing how that is used in the family of harmonic exponential distributions to represent a varied set of densities (Sec. III-B), and finally how these concepts can be combined to perform Bayesian filtering (Sec. III-C).
III-A Harmonic Analysis on Locally Compact Motion Groups
@@
Motion groups are a specific category of Lie groups [18] that serve as a representation of rigid-body motions within the Euclidean space $\mathbb{R}^{N}$. This classification includes other important subgroups such as rotation groups $SO(N)$ and affine transformation groups $SE(N)$. More precisely, these are Lie groups because they satisfy the group axioms (see [26, Chapter 7]) and are also a differentiable manifold. Groups are equipped with a set $G$, alongside a binary operator $\circ$, which preserves the closure property, that is, $\forall g,h\in G$, their composition $g\circ h\in G$.
@@
Conducting Fourier analysis on functions supported on Lie groups necessitates a collection of carefully chosen basis functions. These basis functions are in turn the matrix elements of irreducible unitary representations (IURs). A more concrete definition of IURs is provided below.
A representation denotes a mapping $R:G\rightarrow GL(\mathcal{V})$ of a given group $G$ to the set of linear invertible transformations of a vector space $\mathcal{V}$ such that it respects the group operator, $R(g)R(h)=R(g\circ h)$, $\forall g,h\in G$. Here, $g\circ h$ denotes composition of group elements whereas $R(g)R(h)$ is matrix multiplication given an appropriate basis for $\mathcal{V}.$
A unitary representation, $U$, is a representation that does not affect the norm of the vectors on which the representation acts, that is, $\forall g\in G,\forall v\in\mathcal{V}:||U(g)v||=||v||$. Given a unitary matrix $F$, an irreducible unitary representation (IUR) is a unitary representation that cannot be block diagonalized further into another representation via consecutive transformations of the form $F^{-1}U(g)F$. An IUR is represented as $T^{\lambda}(g)\in\mathbb{R}^{m\times n}$ where $\lambda\in\mathbb{Z}^{+}$ is used to index the matrix elements of different IURs.
@@
IURs are the foundation for performing harmonic analysis on groups. If we have a function, $f:G\rightarrow\mathbb{R}$, that is square integrable, and an IUR of group $G$, $T^{\lambda}(G)$, then we can perform a generalized Fourier transform to compute the frequency components of $f$ as
$\mathcal{F}[f]_{mn}^{\lambda}=\eta_{mn}^{\lambda}=\int_{G}f(g)T_{mn}^{\lambda}(g)^{\dagger}d(g),$ (1)
@@
where $d(g)$ is the bi-invariant volume element of $G$ and $(\cdot)^{\dagger}$ the Hermitian transpose. Note that matrix elements of the Fourier coefficients $\eta^{\lambda}\in\mathbb{R}^{m\times n}$, and IUR $T^{\lambda}(g)\in\mathbb{R}^{m\times n}$ are indexed by their row and column indices $m$ and $n$. Using orthonormality between different matrix elements of $T(g)$ [13], the generalized inverse Fourier transform is
$\mathcal{F}^{-1}[\eta](g)=f(g)$ $=\sum_{\lambda}\sum_{n,m}\eta_{mn}^{\lambda}T_{nm}^{\lambda}(g).$ (2)
@@
For the work presented in this paper, we use the IUR of $SE(2)$ presented in [26, 27]. For $g(x,y,\theta)\in SE(2)$ with $(x,y)\in\mathbb{R}^{2}$ and $\theta\in[0,2\pi]$ parameterizing an element of $SE(2)$, a variable $\psi\in[0,2\pi)$, bi-invariant volume $d\mu\triangleq dx\,dy\,d\theta\,d\psi$ and complex unit $i$, the Fourier coefficient $\eta^{\lambda}_{mn}$ can be computed as
@@
Chirikjian and Kyatkin provide a thorough description of the computation of the fast Fourier transform (FFT) over 2D motion groups in [26, Section 11.2]. Henceforth, we will omit indices and denote the generalized Fourier transform for (locally) compact motion groups as $\eta=\mathcal{F}[f]$.
In this work, the FFT is used to compute the Fourier coefficients from a set of discrete samples of the function on the group. In practice, to avoid an infinite sum or integration, the Fourier analysis needs to be performed on a locally compact subset of a group. For $SE(2)$, the orientation portion, $SO(2)$, is already compact. For translation, we only consider a predefined locally compact subset of $\mathbb{R}^{2}$, as it is presented in [26].
@@ III-B Harmonic Exponential Distributions
where $\eta$ are the parameters, $T(g)$ the IUR of group $G$, $Z_{\eta}$ the normalizing constant, and ()($\cdot$)( ⋅ ) is an appropriate inner product. For example, $\forall g\in SO(2)$ their IURs are $1\times 1$ matrices of the form $T_{00}^{\lambda}(g)=e^{i\lambda g}$, hence, both $T(g)$ and $\eta$ are vectors whose inner product ()($\cdot$)( ⋅ ) is the standard dot product. Note that this is equivalent to computing (2) on $SO(2)$. The parameters $\eta$ and $Z_{\eta}$ can be efficiently computed via Harmonic analysis from $\phi(g\mid\mathbf{\eta})\triangleq\exp(\mathbf{\eta}\cdot T(g))$ as
$\begin{split}\ln\phi(g\mid\eta)&=\eta\cdot T(g)=\mathcal{F}^{-1}[\eta](g),\\ Z_{\eta}&=\int_{G}\phi(g\mid\eta)T^{0}_{00}(g)d(g)=\mathcal{F}[\phi]^{0}_{00},\end{split}$ (4)
with $\mathcal{F}[\phi]^{0}_{00}$ being the zeroth Fourier coefficient and $T^{0}_{00}(g)=1$. Representing the density with Fourier coefficients in log-likelihood space, as opposed to directly representing the density, ensures the probability density function is non-negative for any set of parameters $\eta$, and the normalizing factor $Z_{\eta}$ ensures that it integrates to 1. From this point forward, we use the shorthand notation $p_{a}(g)\triangleq p_{\eta_{a}}(g)$. The harmonic exponential distribution provides an elegant solution for the product of two distributions, achieved through the element-wise addition of their Fourier coefficients:
@@
Figure 2: Relationship between probability $p(g)$, log-likelihood $\text{ln}(p(g))$, and their respective spectral coefficients $\eta$ and $M$. We model the log-likelihood $\text{ln}(p(g))$ using its spectral coefficients $\eta$. To recover the density $p(g)$, we compute the inverse FFT of $\eta$ and exponentiate the result. The spectral coefficients $M$ are computed by taking the FFT of $p(g)$ and are used to compute convolutions.
@@
$=\mathcal{F}^{-1}[\mathcal{F}[p_{b}]\mathcal{F}[p_{a}]](g)$
$=\mathcal{F}^{-1}[M_{b}M_{a}](g),$
with $M_{a}=\int_{G}p_{a}(g)T(g)d(g)=\mathcal{F}[p_{a}]$ being the Fourier coefficients of the probability density function $p_{a}$, instead of a log-likelihood function as in (4). Due to the nonconmutative nature of $SE(N)$, encoded through $T(g)$, the order of the Fourier coefficients is reversed in (6). Observe that $M_{a}$ and $M_{b}$ are composed via matrix products: $M^{\lambda}=M_{b}^{\lambda}M_{a}^{\lambda}$, $\forall\lambda\in\mathbb{Z}^{+}$. A complete proof is presented by Chirikjian and Kyatkin [26, Section 10.4]. The relation between these spaces can be seen in Fig. 2.

### v2

# Please cite this paper as:

@@ Abstract
Bayesian estimation is a vital tool in robotics as it allows systems to update the robot state belief using incomplete information from noisy sensors. To render the state estimation problem tractable, many systems assume that the motion and measurement noise, as well as the state distribution, are unimodal and Gaussian. However, there are numerous scenarios and systems that do not comply with these assumptions. Existing nonparametric filters that are used to model multimodal distributions have drawbacks that limit their ability to represent a diverse set of distributions. This paper introduces a novel approach to nonparametric Bayesian filtering on motion groups, designed to handle multimodal distributions using harmonic exponential distributions. This approach leverages two key insights of harmonic exponential distributions: a) the product of two distributions can be expressed as the element-wise addition of their log-likelihood Fourier coefficients, and b) the convolution of two distributions can be efficiently computed as the tensor product of their Fourier coefficients. These observations enable the development of an efficient and asymptotically exact solution to the Bayes filter up to the band limit of a Fourier transform. We demonstrate our filter’s performance compared with established nonparametric filtering methods across simulated and real-world localization tasks.  [^1] Project page https://montrealrobotics.ca/hef/
@@ I Introduction
Recursive Bayesian filtering is a powerful approach for state estimation in robotics. It allows the fusion of different sensor streams while producing belief estimations over the state that can be used in downstream tasks such as planning, control, and mapping. To render the state estimation problem computationally tractable, certain assumptions are frequently invoked, including the presumption that the state distribution, as well as the motion and measurement noise, are unimodal and Gaussian distributed. However, there are many scenarios where a unimodal Gaussian belief fails to place a high probability (density) over the true state. Estimates involving uncertain data associations [1], propagating errors through rigid-body motions [2], and measurements that suffer from multipath errors [3] and symmetries [4] all result in distributions that are distorted or multimodal.
@@
Nonparametric filtering offers an alternative to Gaussian-based filters, such as particle filters [7, 8, 9] and histogram filters [10, 11, 12]. However, these approaches are not free of limitations. Particle filters are susceptible to particle deprivation [1, Chapter 4], where the correct state is no longer represented within the particle set. Similarly, histogram filters require resource-intensive computations, including convolutions during motion propagation, and are only accurate up to the quantization error when computing probability bins.
This paper draws on seminal ideas in Fourier analysis for optimal filtering on rotation groups [13, 14, 15, 16] and proposes a novel nonparametric, multimodal filtering approach for motion groups, leveraging harmonic exponential distributions [17]. We name this method the harmonic exponential filter (HEF) and it relies on three key benefits of harmonic exponential distributions. First, the prediction step exploits the Fourier convolution theorem, yielding an efficient tensor product instead of a convolution during motion propagation. Second, the update step of the filter becomes an element-wise addition. Finally, the Fourier coefficient representation of the harmonic exponential distribution represents more naturally smoothly-varying density functions when compared to histogram binning. These steps are illustrated in Fig 1.
@@
  • An asymptotically exact Bayesian filter that operates on a compact subset of motion groups, for band-limited prior, motion, and measurement functions.
  • A nonparametric filtering approach for noncommutative motion groups that captures multimodal distributions.
@@
  • An experimental validation of our approach against classical filtering methods.
@@ II Related Work
Various approaches have extended Bayesian filtering beyond the unimodal Gaussian assumptions of the Kalman filter (KF). For instance, the KF has been adapted to operate with a mixture of parameterized distributions instead of a unimodal Gaussian [18]. Gilitschenski et al. [19] adapted a mixture model to use direction statistics on the manifold of rotation groups. However, a key challenge with such approaches is the growth of mixture components with each iteration. Common approaches to overcome this are based on distance metrics to merge mixture components [20] or resampling [21].
The KF has also been extended to exploit symmetries of the system by characterizing the state of the robot using Lie groups [22]. If the state error is modeled in the Lie algebra of motion groups with a Gaussian belief and then mapped back to the group with the exponential map, a non-Gaussian posterior is obtained on the manifold [5, 2]. However, to propagate the error state over time and attain a tractable solution, these methods rely on second-order approximations. The invariant extended Kalman filter (IEKF) [6] also models the state error in the Lie algebra with a Gaussian density, which restricts it to unimodal distributions.
@@
Another family of nonparametric and multimodal filters is the particle filter which represents the posterior belief using a discrete set of particles [7]. These methods can model general distributions, but are known to suffer from particle deprivation, particularly in high-dimensional state spaces [1, Chapter 4]. Research efforts have been devoted to addressing this problem by proposing better sampling schemes [8, 9], and by changing how particles evolve in the prediction step to better match the posterior distribution[24]. In contrast, our method does not use particles to model the posterior belief and does not suffer from particle deprivation.
Kernel density estimates are a prevalent method to represent nonparametric probability density functions. They can be adapted to handle belief propagation through probabilistic graphical models [25], and the product of distributions within a filtering framework [26]. Similarly, learning-based methods have been used to approximate sampling from a non-Gaussian posterior in SLAM by modeling conditional distributions in a Bayes tree with normalizing flows [27].
Harmonic analysis for optimal filtering has been widely studied on rotation groups. Willsky [13] derived state estimation equations for filtering on the circle using infinite Fourier coefficients and proposed finite-dimensional suboptimal implementations [14]. Filtering with “Fourier densities” has since expanded to groups like $SO(N)$ [16]. Similarly to our work, Lo et al. [15] highlighted the advantages of rotational Fourier densities for the update step on $SO(3)$. Fourier densities have also been applied to non-linear filtering, with Brunn et al. [28, 29] formulating Bayesian prediction and update steps on d-dimensional Euclidean spaces. Unlike these methods, our approach extends these ideas beyond compact rotation groups or Euclidean spaces and applies them to motion groups.
Pfaff et al. [30] proposed a method conceptually similar to ours, called Fourier filtering, which was applied to the unit circle and sphere [31, 32]. They also showed their method applied to the problem of multimodal density estimation on the hypertorus, and addressed arbitrary transition densities in the prediction step [33, 34].
Although these approaches resemble ours in that probability densities are represented by Fourier coefficients, our approach extends this concept in several meaningful ways. First, since we model density functions in the log-density space, the product of distributions becomes an element-wise addition instead of the more intricate tensor product needed for Fourier filtering. Second, by representing distributions with the Fourier coefficients derived from the log-density function, the corresponding probability density function is guaranteed to be non-negative [35]. Lastly, our approach can represent densities on compact subsets of motion groups $SE(N)$, particularly any matrix Lie group and not just on a circle or a sphere (the latter of which is not a Lie group).
@@ III Harmonic Exponential Filtering
Fourier analysis provides a way to parameterize well-behaved functions in $\mathbb{R}^{N}$ (i.e. one that is square integrable $\mathcal{L}^{2}$ on $\mathcal{V}\subseteq\mathbb{R}^{N}$ and has a finite number of discontinuities in any finite interval in $\mathbb{R}^{N}$) using an infinite sum of trigonometric basis functions. This approach has proven to be useful for various engineering applications for several reasons. Moreover, the convolution theorem shows that computing integrals in spatial domains can be replaced by an efficient tensor product in the frequency domain. A generalized form of Fourier analysis can be used on compact subsets of motion groups. The topic is presented in depth by Chirikjian and Kyatkin in [36, Chapters 8,10]. This section presents an overview of our approach. First by providing a brief introduction to Fourier analysis on motion groups, particularly $SE(2)$ (Sec. III-A), then showing how it is used in the family of harmonic exponential distributions to represent a varied set of densities (Sec. III-B), and finally how these concepts can be combined to perform Bayesian filtering (Sec. III-C).
III-A Harmonic Analysis on Motion Groups
@@
Motion groups are a specific category of Lie groups [22] that represent rigid-body motions within the Euclidean space $\mathbb{R}^{N}$. This classification includes other important subgroups, such as rotation groups $SO(N)$ and rigid-body-transformation groups $SE(N)$. Particularly, these are Lie groups because they satisfy the group axioms (see [36, Chapter 7]) and are a differentiable manifold. A group consists of a set $G$ and a binary operator $\circ$ that satisfies the closure property; $\forall g,h\in G$, their composition $g\circ h$ also belongs to $G$.
@@
Conducting Fourier analysis on functions supported on Lie groups, such as $SE(2)$, requires a collection of carefully chosen basis functions. These basis functions are in turn the matrix elements of irreducible unitary representations (IURs). A definition of IURs is provided below.
A matrix representation for $SE(2)$, denoted $R^{\lambda}(g)$, is an infinite-dimensional matrix function that satisfies the homomorphism property $R^{\lambda}(g\circ h)=R^{\lambda}(g)R^{\lambda}(h)$, $\forall g,h\in G=SE(2)$, where $\lambda\in\mathbb{R}^{+}$ indexes the representation. Here, $g\circ h$ represents composition of group elements, while $R^{\lambda}(g)R^{\lambda}(h)$ denotes standard matrix multiplication.
A unitary representation $U^{\lambda}_{mn}(g)$, with matrix elements $m,n\in\mathbb{Z}$, is a representation that satisfies $\langle U^{\lambda}(g)\varphi_{1},U^{\lambda}(g)\varphi_{2}\rangle=\langle\varphi_{1},\varphi_{2}\rangle$ for all $m,n\in\mathbb{Z}$. Here, the inner product is defined as $\langle\varphi_{1},\varphi_{2}\rangle=\frac{1}{2\pi}\int_{0}^{2\pi}\varphi_{1}(\psi)^{\mathsf{H}}\varphi_{2}(\psi)d_{\psi}$ where $(\cdot)^{\mathsf{H}}$ denotes the Hermitian transpose and $\varphi_{i}$ is a function defined on $S^{1}$. An irreducible unitary representation (IUR) is a unitary representation that cannot be block-diagonalized further into another representation. A representation can be block-diagonalized if it is equivalent to the direct sum of other representations [36]. An IUR is represented as $T^{\lambda}(g)\in\mathbb{R}^{m\times n}$.
@@
IURs are the foundation for performing harmonic analysis on motion groups. If we have a function, $f:G\rightarrow\mathbb{R}$, that is square integrable, and an IUR of motion group $G$, $T^{\lambda}(G)$, then we can perform a generalized Fourier transform to compute the frequency components of $f$ as
$\mathcal{F}[f]_{mn}^{\lambda}=\int_{G}f(g)T_{mn}^{\lambda}(g)^{\mathsf{H}}d(g)=\eta_{mn}^{\lambda},$ (1)
@@
where $d(g)$ is the bi-invariant volume element of $G$. Matrix elements of the Fourier coefficients $\eta^{\lambda}\in\mathbb{R}^{m\times n}$, and IUR $T^{\lambda}(g)\in\mathbb{R}^{m\times n}$ are indexed by their row and column indices $m$ and $n$. Exploiting the fact that $T(g)$ is a unitary representation, the generalized inverse Fourier transform is
$\mathcal{F}^{-1}[\eta](g)=\sum_{n,m\in\mathbb{Z}}\int_{0}^{\infty}\eta_{mn}^{\lambda}T_{nm}^{\lambda}(g)\lambda d_{\lambda}=f(g),$ (2)
with $\mathcal{F}\left[\mathcal{F}^{-1}[\eta]\right]=\eta$ and $\mathcal{F}^{-1}\left[\mathcal{F}[f]\right]=f$. Chirikjian and Kyatkin provide a proof of the Fourier inversion formula for $SE(2)$ in [36, Section 10.4.2] and offer a detailed explanation of the fast Fourier transform (FFT) over 2D motion groups in [36, Section 11.2].
@@
In this work we use the IUR of $SE(2)$ presented in [36, 37]. For $g(x,y,\theta)\in SE(2)$ with $(x,y)\in\mathbb{R}^{2}$ and $\theta\in[0,2\pi]$ parameterizing an element of $SE(2)$, a variable $\psi\in[0,2\pi)$, bi-invariant volume $d\mu\triangleq dx\,dy\,d\theta\,d\psi$ and complex unit $i$, the Fourier coefficients $\eta^{\lambda}_{mn}$ are computed as
@@
To make the integral in (1) tractable, we assume that $f$ is supported on compact subsets of the group and use a discrete subset of function samples. For $SE(2)$, the orientation component $SO(2)$ is compact, and for translation, we consider a predefined compact subset of $\mathbb{R}^{2}$, as described in [36]. Additionally, we assume $f$ is bandlimited with $\lambda\in(0,\lambda_{\text{max}}]$ and truncate the infinite sums in (2) to $-k\leq m,n\leq k$. We will omit indices and denote the generalized Fourier transform for compact subsets of motion groups as $\eta=\mathcal{F}[f]$.
@@ III-B Harmonic Exponential Distributions
Figure 2: The relationship between probability density $p(g)$, log-density $\text{ln}(p(g))$, and their spectral coefficients $\eta$ and $M$. We model the log-density $\text{ln}(p(g))$ using $\eta$. To recover the $p(g)$, we compute the inverse FFT of $\eta$ and exponentiate the result. The spectral coefficients $M$ are obtained by taking the FFT of $p(g)$ and are used to compute convolutions.
@@
where $\eta$ are the (spectral) parameters, $T(g)$ is the IUR of group $G$, $Z_{\eta}$ is the normalizing constant, and ()($\cdot$)( ⋅ ) is a shorthand notation for $\eta\cdot T(g)\triangleq\mathcal{F}^{-1}[\eta](g)$. Both notations will be used interchangeably to remain consistent with [17]. The parameters $\eta$ and $Z_{\eta}$ can be computed via harmonic analysis from $\phi(g\mid\mathbf{\eta})\triangleq\exp(\mathbf{\eta}\cdot T(g))$ as
$\ln\phi(g\mid\eta)$ $=\eta\cdot T(g)=\mathcal{F}^{-1}[\eta](g),$ (4a)
$Z_{\eta}$ $=\int_{G}\phi(g\mid\eta)T^{0}_{00}(g)\,d(g)=\mathcal{F}[\phi]^{0}_{00}.$ (4b)
with $\mathcal{F}[\phi]^{0}_{00}$ being the zeroth Fourier coefficient and $T^{0}_{00}(g)=1$. Representing the probability density with Fourier coefficients in log-density space, ensures that the probability density function is non-negative for any set of parameters $\eta$, and the normalizing constant $Z_{\eta}$ ensures that it integrates to 1. From this point forward, we use the shorthand notation $p_{a}(g)\triangleq p_{\eta_{a}}(g)$. The harmonic exponential distribution provides an elegant solution for the product of two distributions, achieved through the element-wise addition of their Fourier coefficients:
@@
$=\mathcal{F}^{-1}[\mathcal{F}[p_{b}]\mathcal{F}[p_{a}]](g)=\mathcal{F}^{-1}[M_{b}M_{a}](g)$
with $\mathcal{F}[p_{a}]=\int_{G}p_{a}(g)T(g)d(g)=M_{a}$ being the Fourier coefficients of the density $p_{a}$, instead of a log-density function as in (4a). Due to the nonconmutative nature of $SE(2)$, encoded through $T(g)$, the order of the Fourier coefficients is reversed in (6). Note that $M_{a}$ and $M_{b}$ are composed via matrix products: $M^{\lambda}=M_{b}^{\lambda}M_{a}^{\lambda}$ [36, Section 10.4]. The relation between these spaces is shown in Fig. 2.